SELF- AND AIR-BROADENING OF ^{12}C^{16}O, ^{13}C^{16}O AND ^{12}C^{18}O AT 2.3 μm

V. MALATHY DEVI, D. CHRIS BENNER, The College of William and Mary, Williamsburg, VA 23187; \underline{MARY ANN H. SMITH}, Science Directorate, NASA Langley Research Center, Hampton, VA 23681; ARLAN W. MANTZ, Dept. of Physics, Astronomy and Geophysics, Connecticut College, New London, CT 06320; KEEYOON SUNG and LINDA R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr.,Pasadena, CA 91109.

High resolution (0.005 cm^{−1}) absorption spectra of CO and two of its isotopologues (^{13}CO and C^{18}O) were recorded between 3550 and 5250 cm^{−1} using the Bruker IFS-125HR Fourier transform spectrometer (FTS) located at the Jet Propulsion Laboratory (JPL) and a specially designed and built coolable 20.38 cm long absorption cell^{a} placed within the sample compartment of the FTS. More than 50 spectra of both pure and air-broadened samples of CO, ^{13}CO and C^{18}O were recorded at various temperatures from 150 K to 298 K, with maximum total pressures up to ∼700 Torr. A multispectrum nonlinear least squares spectrum fitting technique^{b} was used to determine the spectral line shape parameters including speed dependence, Lorentz halfwidth coefficients, pressure-induced shift coefficients, and off-diagonal relaxation matrix element coefficients for line mixing. These line shape parameters were obtained for both self- and airbroadening, and temperature dependences of these parameters were determined where possible. As previously done in studies of CO_{2},^{c} rather than retrieving individual line positions and intensities, we constrained them to their theoretical relationships, including Herman-Wallis terms, determining only the band intensities and rovibrational constants. The results are discussed and compared with values reported in the literature.^{d}

aK. Sung, A. W. Mantz, M. A. H. Smith, L. R. Brown, T. J. Crawford, V. Malathy Devi and D. C. Benner, JMS \underline{262} (2010)

122-134.

bD. C. Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith and D. A. Atkins, JQSRT \underline{53} (1995) 705-721.

cV. Malathy Devi, D. C. Benner, L. R. Brown, C. E. Miller and R. A. Toth, JMS \underline{242} (2007) 90-117.

dResearch described in this paper was performed at Connecticut College, the College of William and Mary, NASA Langley Research Center and the Jet Propulsion Laboratory, California Institute of Technology, under contracts and cooperative agreements with the National Aeronautics and Space Administration.